OpenStudy (anonymous):

what is equivalent to log36 9

OpenStudy (anonymous):

the base is 36 ?

OpenStudy (anonymous):

yes logV36 ^9

OpenStudy (anonymous):

the 9 isnt an exponent.

OpenStudy (anonymous):

\[\log _{36}9\]\[\frac{\log3^2}{\log6^2}=\frac{2\log(3)}{2\log(2)+2\log(3)}\]

OpenStudy (anonymous):

=0.613147192

OpenStudy (anonymous):

\[\frac{\log 9}{ \log 36} =0.613147192\]